# Flow Straighter and Faster: Efficient One-Step Generative Modeling via MeanFlow on Rectified Trajectories

### v1

@@ B.4 Training with Guidance
MeanFlow [11] further introduced an improved CFG method that mixes conditional and unconditional mean-velocity predictions:
@@
Empirically, we found that directly training MeanFlow on the CFG field results in instability, consistent with observations in Hu et al. [19]. To mitigate this, we use a simple two-stage strategy: first train $u_{\theta}$ on the unconditional flow, then on the CFG-modified flow. Usually, allocating half of the total training budget to each stage provides a good balance between stability and final quality.
@@
We also found that sampling a random CFG scale $\omega^{\prime}$ from a uniform distribution (rather than fixing it) gives better results. Large values of $\kappa$ is also important for stable training. In practice, we sample $\omega^{\prime}$ from a uniform distribution, then set $\kappa=\max(1.0,\omega^{\prime}-1)$, and finally compute the corresponding value for $\omega=\frac{\omega^{\prime}}{1-\kappa}$.
B.5 2D Toy Example Setting
We consider a controlled two-dimensional transport task designed to highlight curvature effects and the behavior of different flow-learning methods. Both the source and target distributions are mixtures of two Gaussians. The source distribution is a balanced mixture with equal component weights, while the target distribution is an imbalanced mixture with weights $(0.6,0.4)$. Component means are placed symmetrically on the left (source) and right (target) sides of the plane, and all components use identity covariance. This setup ensures that samples associated with the lighter target mode must traverse longer paths, creating the geometric imbalance that induces curved trajectories.
All velocity and MeanFlow models are implemented as small multilayer perceptrons (MLPs) with identical architectures across methods. Training is performed for a fixed budget of 20k steps using Adam with learning rate $10^{-3}$ and batch size 1024.
We compare three approaches under this shared budget:
  1. 1.
    2-rectified Flow: two successive reflow steps applied to independently sampled couplings, each trained for 10k iterations.
  2. 2.
    MeanFlow: trained for the full 20k iterations directly on independent couplings.
  3. 3.
    Re-Meanflow (ours): a velocity model is first trained for 10k iterations to obtain a 1-rectified flow; a MeanFlow model is then trained for another 10k iterations on the resulting rectified couplings.
Figure 3 visualizes the learned trajectories and one-step Euler sampling results for all three settings. As discussed in the main text, 2-rectified flow retains noticeable outliers due to residual curvature, and MeanFlow trained on independent couplings fails to converge within the budget. In contrast, Re-Meanflow combines moderate rectification with stable MeanFlow training to achieve accurate one-step generation.
Appendix C More Experiment Details and Results
In this section, we provide additional experimental details, computational analyses, and extended results. We describe how we estimate FLOPs and GPU-hours for all compared methods, and we provide additional distance-error correlation for ImageNet $64^{2}$ and $256^{2}$.
C.1 Computation Estimation of Each Method
FLOPS Estimation.
@@ C.1 Computation Estimation of Each Method
In Fig 7, we report the efficiency in terms of estimated exaFLOPs (Eflops). To ensure comparability, we estimate total training and sampling compute for each method based on their reported FLOPs per forward pass. Specifically, we use the following assumptions:
@@
  • The FLOPs of a forward pass are reported by prior works (e.g., EDM[21]: $100$ GFLOPs, EDM2-S[23]: $102$ GFLOPs, SiT-XL[31]: $118.64$ GFLOPs).
@@
    $\text{Total Train FLOPs}=(\#\text{iters})\times(\text{batch size})\times(\text{forward}+\text{backward})\times(\text{GFLOPs per fwd}).$
@@
    $\text{Total Sample FLOPs}=(\#\text{samples})\times(\#\text{steps})\times(\text{forward passes per step})\times(\text{GFLOPs per fwd}).$
@@
    $\underbrace{5\times 10^{6}}_{\text{\#samples}}\times\underbrace{63}_{\text{steps}}\times\underbrace{2}_{\text{fwd/step (auto-guidance)}}\times\underbrace{102}_{\text{GFLOPs/fwd}}\approx 6.43\times 10^{10}\;\;\text{GFLOPs}\;\;\approx\;\;64\;\text{Eflops}.$
@@
    $\underbrace{50{,}000}_{\text{iters}}\times\underbrace{128}_{\text{batch}}\times\underbrace{(1+4)}_{\text{fwd+back}}\times\underbrace{102}_{\text{GFLOPs/fwd}}\;\;+\;\underbrace{50{,}000}_{\text{iters}}\times\underbrace{128}_{\text{batch}}\times\underbrace{(3+4)}_{\text{fwd+back}}\times\underbrace{102}_{\text{GFLOPs/fwd}}\approx 8\;\text{Eflops}.$
  • For training, the total compute is computed as:
    $\text{Total Train FLOPs}=(\#\text{iters})\times(\text{batch size})\times(\text{forward}+\text{backward})\times(\text{GFLOPs per fwd}).$
  • For sampling in the reflow process, the total compute is computed as:
    $\text{Total Sample FLOPs}=(\#\text{samples})\times(\#\text{steps})\times(\text{forward passes per step})\times(\text{GFLOPs per fwd}).$
    $\underbrace{5\times 10^{6}}_{\text{\#samples}}\times\underbrace{63}_{\text{steps}}\times\underbrace{2}_{\text{fwd/step (auto-guidance)}}\times\underbrace{102}_{\text{GFLOPs/fwd}}\approx 6.43\times 10^{10}\;\;\text{GFLOPs}\;\;\approx\;\;64\;\text{Eflops}.$
    $\underbrace{50{,}000}_{\text{iters}}\times\underbrace{128}_{\text{batch}}\times\underbrace{(1+4)}_{\text{fwd+back}}\times\underbrace{102}_{\text{GFLOPs/fwd}}\;\;+\;\underbrace{50{,}000}_{\text{iters}}\times\underbrace{128}_{\text{batch}}\times\underbrace{(3+4)}_{\text{fwd+back}}\times\underbrace{102}_{\text{GFLOPs/fwd}}\approx 8\;\text{Eflops}.$
  • Example: AYF [37]. AYF does not require sampling, so only the training compute is considered:
@@
GPU Hours Estimation.
In Fig.7, we also estimate the total GPU hours for AYF[37] and 2-rectified flow++ [24]. For all methods, including ours, we follow the standard convention of computing GPU hours as
@@
C.2 Distance-Error Correlation at Other Resolutions
In Fig. 9, we replicate the distance-error correlation analysis of Fig. 4 on ImageNet $256^{2}$ (Left) and $64^{2}$ (Right), respectively. In both settings, we observe the same phenomenon: couplings with large data-noise $\ell_{2}$ distances form a pronounced high-error tail of the flow regression objective. This mirrors our ImageNet $512^{2}$ results.
Figure 9: Distance-Error Correlation. Histogram of data-noise $\ell_{2}$ distances on ImageNet $256^{2}$ (Left) and ImageNet $64^{2}$ (Right), analogous to Fig. 4. Similar high-distance, high-error tails (90th percentile marked) are observed.
Appendix D Algorithm for Re-Meanflow
Algorithm 1 Re-Meanflow Training
1:dataset $\mathcal{D}$; prior $p_{\mathbf{z}}$; time distributions $p_{t}$ and $p_{r,t}$; number of rectified-flow iterations $T_{\text{flow}}$; number of meanflow iterations $T_{\text{MF}}$; number of coupling samples $N_{\text{pairs}}$; truncation ratio $k\%$.
2:
3:Stage 1: Train 1-rectified Flow on Independent Couplings
4:for $\text{iter}=1$ to $T_{\text{flow}}$ do
5:  Sample $\mathbf{x}\sim\mathcal{D}$, $\mathbf{z}\sim p_{\mathbf{z}}$, $t\sim p_{t}$
6:  $\mathbf{z}_{t}\leftarrow(1-t)\mathbf{x}+t\mathbf{z}$
7:  $\dot{\mathbf{z}}_{t}\leftarrow\mathbf{z}-\mathbf{x}$
8:  Minimize $\|\dot{\mathbf{z}}_{t}-v_{\theta}(\mathbf{z}_{t},t)\|_{2}^{2}$ w.r.t. $\theta$
9:end for
10:Freeze the resulting 1-rectified velocity field $v^{1}_{\theta}$
11:
12:Stage 2: Build Rectified Couplings + Distance Truncation
13:for $n=1$ to $N_{\text{pairs}}$ do
14:  Sample $\mathbf{x}\sim\mathcal{D}$
15:  Solve backward ODE with $v^{1}_{\theta}$ from $t=1$ to $t=0$ to obtain $\mathbf{z}$
16:  $d\leftarrow\|\mathbf{x}-\mathbf{z}\|_{2}$
17:  Store $(\mathbf{x},\mathbf{z},d)$
18:end for
19:Let $q$ be the (100k)(100-k)th percentile of distances $\{d_{i}\}$
@@ Appendix D Algorithm for Re-Meanflow
21:
22:Stage 3: Train MeanFlow on Rectified, Truncated Couplings
23:for $\text{iter}=1$ to $T_{\text{MF}}$ do
24:  Sample $(\mathbf{x},\mathbf{z})\sim\mathcal{D}_{\text{rect}}$, $(r,t)\sim p_{r,t}$
25:  $\mathbf{z}_{t}\leftarrow(1-t)\mathbf{x}+t\mathbf{z}$
26:  Compute $g=(t-r)\frac{d}{dt}u_{\phi}(\mathbf{z}_{t},r,t)$ $\triangleright$ JVP
27:  $u_{\text{tgt}}\leftarrow v^{1}_{\theta}(\mathbf{z}_{t},t)-g$
28:  Minimize $\|u_{\phi}(\mathbf{z}_{t},r,t)-\operatorname{sg}(u_{\text{tgt}})\|_{2}^{2}$ w.r.t. $\phi$
29:end for
30:trained meanflow model $u_{\phi}$
Appendix E More Qualitative Results

### v2

@@ 0.B.4 Training with Guidance
MeanFlow [geng2025mean] further introduced an improved CFG method that mixes conditional and unconditional mean-velocity predictions:
@@
Empirically, we found that directly training MeanFlow on the CFG field is unstable, consistent with observations in [hu2025cmt]. To mitigate this, we use a simple two-stage strategy: first train $u_{\theta}$ on the unconditional flow, then on the CFG-modified flow. Usually, allocating half of the total training budget to each stage provides a good balance between stability and final quality.
@@
We also found that sampling a random CFG scale $\omega^{\prime}$ from a uniform distribution (rather than fixing it) gives better results. Large values of $\kappa$ are also important for stable training. In practice, we sample $\omega^{\prime}$ from a uniform distribution, then set $\kappa=\max(1.0,\omega^{\prime}-1)$, and finally compute the corresponding value for $\omega=\frac{\omega^{\prime}}{1-\kappa}$.
Figure 9: Loss landscape of $\mathcal{L}_{MF}$ on a PCA plane of $\mathbf{z}_{t}$ for ImageNet $64^{2}$ (a) and $256^{2}$ (b). We evaluate $\mathcal{L}_{MF}$ on a 2D PCA grid with $t=0.7$ and $r\in\{0.7,0.5,0.3,0.1\}$ (left to right). MeanFlow (top) becomes increasingly sharp and irregular as trt-r grows, whereas Re-MeanFlow (bottom) remains substantially smoother and better-conditioned, supporting more stable and efficient training.
Appendix 0.C More Experiment Details and Results
In this section, we provide additional experimental details, computational analyses, and extended results. We first present loss-landscape visualizations at $64^{2}$ and $256^{2}$, then describe our protocol for estimating FLOPs and GPU-hours across all methods. Finally, we report additional convergence results that further support the synergy between trajectory rectification and mean-velocity modeling.
0.C.1 Loss Landscape
In Fig. 7, we visualize the loss landscape of $\mathcal{L}_{MF}$ with respect to the input $\mathbf{z}_{t}$ for MeanFlow and Re-MeanFlow across different $(t,r)$ pairs. Here we report the corresponding results at additional resolutions in Fig. 9, following the same protocol.
Concretely, we fix $t=0.7$ and sweep $r$, and compute PCA directions by collecting 2020\,k samples of $\mathbf{z}_{t}$ (equivalently, $\mathbf{x}_{t}$ in our implementation). For each resolution, we then evaluate $\mathcal{L}_{MF}$ on a 2D grid spanned by the top two PCA directions using a minibatch of samples. Collecting the data and generating all surfaces for a single resolution takes approximately $3$ hours on one A100 GPU. Overall, consistent with Fig. 7, Re-MeanFlow exhibits a substantially smoother and better-conditioned loss landscape, which translates into more efficient and stable training.
0.C.2 Computation Estimation of Each Method
0.C.2.1 FLOPS Estimation.
@@ 0.C.2 Computation Estimation of Each Method
In Fig. 6 (left), we report efficiency in terms of estimated exaFLOPs (EFLOPs). To ensure comparability, we estimate total training and sampling compute for each method based on their reported FLOPs per forward pass. Specifically, we use the following assumptions:
@@
  • The FLOPs of a forward pass are reported by prior works (e.g., EDM[karras2022elucidating]: $100$ GFLOPs, EDM2-S[karras2024analyzing]: $102$ GFLOPs, SiT-XL[ma2024sit]: $118.64$ GFLOPs).
@@
    $\text{Total Train FLOPs}=$ $(\#\text{iters})\times(\text{batch size})\times(\text{forward}+\text{backward})$
    $\times(\text{GFLOPs per fwd}).$
@@
    $\text{Total Sample FLOPs}=$ $(\#\text{samples})\times(\#\text{steps})\times(\text{forward passes per step})$
    $\times(\text{GFLOPs per fwd}).$
@@
    $\underbrace{5\times 10^{6}}_{\text{\#samples}}\times\underbrace{63}_{\text{steps}}\times\underbrace{2}_{\text{fwd/step (auto-guidance)}}\times\underbrace{102}_{\text{GFLOPs/fwd}}\approx\;\;64\;\text{Eflops}.$
@@
    $\underbrace{50{,}000}_{\text{iters}}\times\underbrace{128}_{\text{batch}}\times\underbrace{(1+4)}_{\text{fwd+back}}\times\underbrace{102}_{\text{GFLOPs/fwd}}\;\;$
    $+\;\underbrace{50{,}000}_{\text{iters}}\times\underbrace{128}_{\text{batch}}\times\underbrace{(3+4)}_{\text{fwd+back}}\times\underbrace{102}_{\text{GFLOPs/fwd}}\approx 8\;\text{Eflops}.$
  • Example: AYF [sabour2025align]. AYF does not require sampling, so only the training computation is considered:
@@
0.C.2.2 GPU Hours Estimation.
In Fig.6, we also estimate the total GPU hours for AYF[sabour2025align] and 2-rectified flow++ [lee2024improving]. For all methods, including ours, we follow the standard convention of computing GPU hours as
@@
0.C.3 Additional Convergence Comparison with Other Related Methods
In Fig. 1c, we compare the convergence of Re-MeanFlow with MeanFlow [geng2025mean]. Here, we provide additional convergence results against closely related baselines.
0.C.3.1 2-rectified flow++.
We compare against 2-rectified flow++ [lee2024improving], which improves Rectified Flow but still models the instantaneous velocity on rectified couplings, in contrast to Re-MeanFlow, which models mean-velocity. We observe that one-step sample quality under 2-rectified flow++ improves slowly. We attribute this to the sensitivity of one-step Euler updates to residual curvature: accurate one-step sampling with instantaneous velocity effectively requires near-perfect straightness, since even mild curvature can cause overshooting. Because rectified trajectories are not perfectly linear in practice, this sensitivity compounds and slows convergence. This observation aligns with our efficiency comparison in Sec. 4.2: relative to 2-rectified flow++, Re-MeanFlow achieves a $33.4\%$ lower FID while being $26\times$ faster.
0.C.3.2 MeanFlow with Mini-batch OT couplings.
We also train MeanFlow using Mini-batch OT[tong2023improving] couplings, obtained by locally solving an OT matching within each mini-batch (via an OT server) to pair sampled noises and data points. Results on ImageNet $512^{2}$ are shown in Fig. 10a. While mini-batch OT slightly improves over standard MeanFlow, its convergence behavior remains similar and still lags substantially behind Re-MeanFlow. This is consistent with our discussion in Sec. 2.2: Mini-batch OT provides only local transport structure and does not guarantee globally straight trajectories.
Overall, Re-MeanFlow converges substantially faster than these closely related alternatives, highlighting the synergy of modeling mean-velocity on rectified trajectories.
Figure 11: Curvature estimation of generative trajectories on ImageNet $64^{2}$ (a) and $256^{2}$ (b). Using the proxy in Sec. 3.3, we visualize $\mathrm{Curv}(r,t)$ over the $(r,t)$-plane. Rectified couplings used by Re-MeanFlow induce substantially straighter trajectories (lower curvature) than the independent couplings used by MeanFlow.
0.C.4 Curvature Estimation
In Fig. 2, we visualize the curvature proxy on ImageNet $512^{2}$. Here we provide the corresponding curvature estimates at additional resolutions in Fig. 11. The same trend holds across settings: compared to independent couplings, the rectified couplings used to train Re-MeanFlow consistently induce substantially lower curvature over the $(r,t)$-plane.
Figure 12: A 2D Toy Example. We consider a controlled 2D setup where a flow model transports a two-component Gaussian mixture on the left to a mixture on the right. Panels (a-c): (a) Linear interpolation of independently sampled couplings $p_{\mathbf{x}}\times p_{\mathbf{z}}$, which serves as the training signal for the first velocity model. (b) The resulting 1-rectified flow learned from these independent couplings; the learned velocity field remains noticeably curved. (c) Using the velocity field from (b), we generate a new set of couplings and train a second velocity model on their linear interpolations, yielding the 2-rectified flow. Panel (d): Due to imperfect straightening, one-step Euler sampling on the 2-rectified flow still yields noticeable outliers. Panel (e): MeanFlow trained directly on independent couplings fails to converge within the training budget because high-variance conditional velocities destabilize learning. Panel (f): Re-MeanFlow combines trajectory rectification with mean-velocity modeling, eliminating most outliers and achieving more accurate one-step generation.
Figure 13: Distance-based truncation removes high-curvature couplings and improves one-step behavior in a 2D toy task. (a) Independent couplings visualized by linear interpolation between sampled endpoints. (b) Rectified couplings obtained by multi-step sampling with 1-rectified flow, where a small set of long-distance “diagonal” pairings induces high-curvature trajectories. (c) Retained trajectories after truncating the top $10\%$ couplings ranked by endpoint distance $\|\mathbf{x}-\mathbf{z}\|_{2}$ (bottom $90\%$). (d) Removed trajectories (top $10\%$) discarded by truncation. (e) Re-MeanFlow trained on the untruncated rectified couplings in (b), exhibiting residual outliers. (f) Re-MeanFlow trained on the truncated couplings in (c), yielding much cleaner one-step transport with outliers largely eliminated.
Appendix 0.D 2D Toy Example
To visualize the synergy of Re-MeanFlow relative to MeanFlow[geng2025mean] and Rectified Flow [liu2022flow], we present a 2D toy transport problem in Fig. 12. The task maps a two-mode Gaussian mixture to another two-mode mixture. We compare one-step generation under a fixed budget of 20k training iterations for three approaches: (i) 2-rectified flow, (ii) MeanFlow, and (iii) Re-MeanFlow (ours), which trains a velocity model for 10k iterations to obtain a 1-rectified coupling, followed by 10k iterations of MeanFlow on the resulting couplings.
Fig. 12d: 2-rectified flow.
Even after one rectification step, some samples traverse disproportionately long distances toward the denser mode (e.g., due to approximation error or mild mode imbalance). As a result, nearby noise samples can map to widely separated data points, yielding a large effective $L$ in Eq. 10 and inducing early-time curvature near $t\!\approx\!0$. The pronounced one-step deviations in Fig. 12d illustrate this failure mode: one-step Euler updates using instantaneous velocity can overshoot when trajectories are not perfectly straight. Similar geometric imbalances can arise in real datasets, motivating robustness to residual curvature.
Fig. 12e: MeanFlow.
MeanFlow trained directly on independent couplings struggles to learn a coherent transport map within the same budget, resulting in poor one-step generation.
Fig. 12f: Re-MeanFlow (ours).
Re-MeanFlow achieves more accurate one-step generation with substantially fewer invalid samples. This illustrates the complementary roles of the two components: rectification reduces curvature enough to make MeanFlow optimization efficient, while mean-velocity modeling removes the requirement of perfectly straight trajectories.
0.D.0.1 Distance-based Truncation
To illustrate the effect of distance-based truncation, we perform an additional toy experiment shown in Fig. 13. We discard the top $10\%$ of couplings ranked by endpoint distance $\|\mathbf{x}-\mathbf{z}\|_{2}$. Fig. 13d visualizes the retained (truncated) trajectories. While this criterion does not guarantee removing all high-curvature paths, it eliminates most of the “diagonal” couplings that exhibit the highest curvature in the toy setting (as analyzed above). Consequently, Re-MeanFlow trained on the truncated couplings (Fig. 13f) achieves even cleaner one-step generation, removing nearly all outliers compared to training on the untruncated couplings (Fig. 13e).
Appendix 0.E More Qualitative Results